OpenStudy (adamk):

When rolling a 6-sided die, what's the probability of having to roll 6 times before you get a 4? I'm trying to use binomial and geometric functions here.

satellite73 (satellite73):

at least six times?

OpenStudy (mhchen):

Since this doesn't have a set number of trials, it's a geometric distribution.

satellite73 (satellite73):

or exactly six times?

OpenStudy (adamk):

@mhchen Yeah I figured that much out.

OpenStudy (adamk):

@satellite73 Exactly 6 times

OpenStudy (shawn):

######4 is what you want

OpenStudy (shawn):

where # means a number besides 4

OpenStudy (mrs.ambrose614):

Ok first you have to fraction out the possibility of rolling on that number

OpenStudy (mhchen):

So the success is 1/6 ofc And for the geometric formula at P(X=6) is \[p(1-p)^{n-1}\] Since it's 5 failures and 1 success on the 5th trial. So then you plug in the numbers \[(\frac{1}{6})(1-\frac{1}{6})^{6-1}\]

OpenStudy (mhchen):

Sorry I meant *6th trial

OpenStudy (adamk):

Geocdf?

OpenStudy (adamk):

I got 0.067

satellite73 (satellite73):

interesting, because i never know the names of these distributions not 6, not 6, not 6, not 6, not 6, 6\[\left(\frac{5}{6}\right)^5\times \frac{1}{6}\]

OpenStudy (shawn):

(5/6)^6 * (1/6) = 0.056

satellite73 (satellite73):

it keeps you from writing silly stuff like \[(1-\frac{1}{6})\]

OpenStudy (mhchen):

Yeah, same here, about 0.067ish

OpenStudy (mhchen):

I'm not really a 'logical thinker' :b I rely more on formulas

OpenStudy (adamk):

Yeah same here. That's why I was wanting to know what goemetric function it uses. I don't need to know the formulas.